# Optimizing the potential of KM3NeT in detecting core-collapse supernovae

### v1

@@ 5.2 Systematic uncertainties
The main source of uncertainties on the background rates is the extrapolation of the fraction of events surviving the triggered event veto, or “veto acceptance”, when considering future detector configurations. Radioactivity backgrounds are not affected by the triggered event veto, whereas atmospheric muon backgrounds are strongly suppressed in a manner that is dependent on the detector geometry [van2021km3net]. Consequently, for a given event sample, the veto acceptance is determined by two quantities: (1) the relative fraction of radioactivity events in the sample before the veto, which is obtained through the data-driven method described in Section 4, and (2) the fraction of atmospheric muon events that are found to survive the veto, evaluated using simulations. Following the procedure from [van2021km3net], uncertainties on the veto acceptance are estimated by comparing data and simulations for the ORCA15 and ARCA21 configurations for each value of the multiplicity and the BDT discriminant. To avoid underestimating the systematic uncertainties, simulations are corrected before the triggered event veto is applied. The veto acceptance as a function of the multiplicity is presented in Figure 7 for both ORCA and ARCA, showing that discrepancies between data and simulations and the resulting systematic uncertainties are well under control.
@@
Systematic uncertainties on the neutrino interaction cross section, water optical properties, and PMT characteristics are also considered both for the cut optimization and when computing sensitivities to CCSN models. The uncertainty values from [van2021km3net] are used for the cut optimization, and these values are re-evaluated when computing the sensitivities.
@@ 6 KM3NeT sensitivity to CCSN neutrinos
A key objective of this analysis is to position KM3NeT as a strong contributor to the Supernova Early Warning System (SNEWS) [al2021snews], which combines the information received from multiple detectors to reliably identify a CCSN (with a false alarm rate of one per century), locate the supernova using triangulation, and notify telescopes for prompt electromagnetic follow-up. The procedure developed in this paper is designed to identify potential CCSN neutrino signals and issue rapid alerts to this global network of observatories and detectors, following the realtime analysis procedure described in [aiello2022implementation]. In this section, the sensitivity of the new CCSN analysis based on BDTs, incorporating the selection steps and systematic uncertainty evaluation procedure described above, is estimated for different detector configurations and CCSN models. The number of selected single-DOM events is computed within a $0.5\text{\,}\mathrm{s}$ time window, corresponding to the high-rate neutrino emission phase predicted by most models, and compared to the background expectation. As in the previous analysis [van2021km3net, aiello2022implementation], the selection steps are optimized to maximize the distance up to which a CCSN with a neutrino flux corresponding to the $11~M_{\odot}$ model from [tamborra2014neutrino] can be detected at $5\sigma$.
@@ 6.1 Selection procedure
In the previous CCSN analysis [van2021km3net], the optimal multiplicity range was defined using a simple analytical approximation of the signal detection significance, neglecting systematic uncertainties. In this analysis, the signal detection significance is evaluated using the Rolke method [rolke2005limits], which accounts for systematic uncertainties.
@@
Four detector configurations have been considered to optimize the BDT and multiplicity cuts for a CCSN spectrum corresponding to the $11~M_{\odot}$ model considered in [van2021km3net]: ARCA29, ORCA24, ARCA230 and ORCA115. The former represent the first ARCA and ORCA configurations for which the method presented here was used to search for CCSNe in real time. The latter represent the final ARCA and ORCA configurations. In each case, the background is modeled using data from the ARCA21 and ORCA15 configurations, respectively, and the signal is simulated using the average PMT efficiencies associated with these detectors. Since the typical $f_{A}$ values for ORCA15 range between $97$ and $98\%$, the associated background rates have been rescaled to model the $f_{A}>99\%$ scenario. With this choice of detector configurations, for ARCA29 and ORCA24, background rates can be rescaled linearly to account for the increase in detector size, as the variation of the background rate per DOM between the different configurations is negligible. For ARCA230 and ORCA115, an additional step is required to account for the reduced veto acceptance for background muons resulting from the increase in detector size. For this, the background evaluation method described in [van2021km3net] is used: the background rates are evaluated from data before applying the triggered event veto, rescaled to match the increase in the number of DOMs, and then multiplied by the veto acceptances obtained from simulations (shown in Figure 7).
@@
The numbers of selected CCSN candidates after the optimization procedure are shown in the top panel of Figure 8 for the ORCA24 and ARCA29 detector configurations and for ORCA115-ARCA230. The optimal multiplicity ranges for the ORCA24-ARCA29 and ORCA115-ARCA230 detector configurations are 6116-11 and 7107-10, respectively, close to the range found in [van2021km3net]. As in the previous KM3NeT CCSN analysis, ORCA backgrounds are less dominant thanks to ORCA’s higher atmospheric muon identification capabilities at low energies. For the signal, the $11~M_{\odot}$, $27~M_{\odot}$, and $40~M_{\odot}$ models used in [van2021km3net] are shown for a CCSN at $10\text{\,}\mathrm{kpc}$. For all three models, the signal is dominating over the background, indicating that even with a partial configuration, KM3NeT is sensitive to most CCSN candidates in the dense Galactic bulge. To assess the impact of the BDT selection, the signal and background distributions obtained before this selection are shown in the bottom panel of Figure 8. The background rates differ from [van2021km3net] as they have been evaluated using the method described in Section B.2 but the resulting impact on KM3NeT’s detection horizon is negligible. Applying BDT selection criteria significantly improves the signal-to-background ratio, in particular by improving the muon rejection rate at ARCA.
@@
Figure 8: Numbers of selected CCSN candidates for signal and background using the BDT-based method presented in this article (top) and the method from [van2021km3net] (bottom), as a function of the multiplicity for ORCA24-ARCA29 (left) and ORCA115-ARCA230 (right). The multiplicity range shown for each detector configuration and selection procedure is the one providing the optimal sensitivity, as discussed in Section 6.1. For the signal, the $11~M_{\odot}$, $27~M_{\odot}$, and $40~M_{\odot}$ models used in [van2021km3net] are shown. The light-blue upward-pointing and dark-blue downward-pointing triangles represent the ORCA and ARCA backgrounds, respectively. The dark blue histogram is the sum of the two backgrounds.
@@ 6.2 Sensitivities to CCSN models
Table 1: Systematic uncertainties of the analysis. The third column indicates the resulting systematic uncertainties on the signal and background rates denoted as (S) and (B), respectively. The evaluation of the uncertainties on the muon veto acceptance and single-DOM signal event modeling is described in Section 5.2. These uncertainties are quoted here for the optimal BDT cuts corresponding to the recent and final detector configurations. The monthly variability of the instrumentation efficiency [aiello2022implementation] has been evaluated over a period of 11 year. The variations and uncertainties on the water transparency, the finite generation volume and the inverse beta decay (IBD) and elastic scattering (ES) cross sections are taken from [van2021km3net].
@@
    In [van2021km3net], the source of the corresponding uncertainty was considered to be the fraction of active PMTs. “Instrumentation efficiency variability” is a more accurate terminology which was introduced in a subsequent study [aiello2022implementation].
Sensitivities are evaluated for the CCSN signals in the case of the $11~M_{\odot}$, $27~M_{\odot}$, and $40~M_{\odot}$ progenitor mass models and are shown in Figure 9, as a function of the distance to the CCSN, for the ORCA24-ARCA29 (left) and ORCA115-ARCA230 (right) detector configurations. Introducing the BDT selection leads to a $25$ to $30\%$ increase of the $5\sigma$ distance horizon as compared to the one obtained with the method from [van2021km3net]. For the $11~M_{\odot}$ CCSN model, associated with the lowest neutrino flux, KM3NeT’s recent partial configurations have been sensitive to CCSNe up to a distance of $13\text{\,}\mathrm{kpc}$ and the detector will achieve full Galactic sensitivity when completed. The current reach of KM3NeT allows probing at least $79\%$ of Galactic CCSN candidates [Adams:2013ana], $46\%$ more than with the previous analysis strategy [van2021km3net]. The results highlight both KM3NeT’s potential for characterizing MeV neutrino bursts and the interest of investigating the structure of single-DOM hit coincidences to study low-energy signals.
@@
Figure 9: Sensitivity, in units of $\sigma$, as a function of the CCSN distance to Earth for the $11~M_{\odot}$ [tamborra2014neutrino], $27~M_{\odot}$ [tamborra2014neutrino], and $40~M_{\odot}$ [walk2020neutrino] models already considered in the previous KM3NeT CCSN search [van2021km3net]. Left: sensitivities for the ORCA24-ARCA29 detector configuration. Right: projected sensitivities for the completed ORCA115-ARCA230 detector. The sensitivities obtained with the selection from  [van2021km3net] are indicated by the dotted lines. For ORCA115-ARCA230, these detection horizons agree with the ones from [van2021km3net] despite the different background estimation method.
Figure 10: Sensitivity, in units of $\sigma$, as a function of the CCSN distance to Earth for the $11~M_{\odot}$ [tamborra2014neutrino], $27~M_{\odot}$ [tamborra2014neutrino], and $40~M_{\odot}$ [walk2020neutrino] models already considered in the previous KM3NeT CCSN search [van2021km3net]. In this figure, the performances of ARCA and ORCA are evaluated separately. The sensitivities for the ORCA24 (solid lines) and ARCA29 (dashed lines) detector configurations are indicated in the left panel, and the projected sensitivities for the ORCA115 (solid lines) and ARCA230 (dashed lines) detectors are indicated in the right panel.
@@ 6.3 Impact of variations in data-taking conditions
The sensitivities presented above have been evaluated assuming a fraction of active PMTs $f_{A}>99\%$. While this value is now achieved for more than $99\%$ of ARCA’s data-taking period, $f_{A}$ can reach values as low as $50\%$ at ORCA, where bioluminescence levels are typically higher. To account for these variations, the parameterization of the background rate dependence on $f_{A}$ proposed in [aiello2022implementation] is adapted to the BDT-based single-DOM event selection procedure. The rate of single-DOM background events as a function of $f_{A}$ is shown in Figure 11 for ORCA18. For more than $95\%$ of ORCA18’s data-taking period, background rates can be successfully predicted using a parameterization obtained for the ORCA6 configuration. This predictability enables a reliable estimation of the expected background level when searching for CCSNe in real time, even when the bioluminescence activity is high. For ORCA, during $68\%$ of the data-taking time, $f_{A}$ will range from $91\%$ to $99\%$, resulting in variation of ORCA’s CCSN detection horizon by only $8\%$. The variation of the combined ORCA+ARCA detection horizon is expected to be even smaller due to the lower bioluminescence activity (hence typically larger $f_{A}$) at ARCA. Hence, KM3NeT’s CCSN identification capabilities are robust even in the highly dynamic environment of the Mediterranean Sea.
@@ 7 Conclusions
In this paper, a novel approach is proposed to improve the sensitivity of KM3NeT to CCSN neutrinos, by selecting single-DOM hit clusters based on their topology. A multi-variate analysis using BDTs has demonstrated that most of the hit information in a single-DOM event can be encapsulated into a small set of observables. With these observables, the ability to identify CCSN neutrino signals and separate them from radioactivity and atmospheric muon signatures, thereby allowing the estimation of the corresponding systematic uncertainties, is enhanced. Consequently, KM3NeT’s CCSN detection horizon is extended by $30\%$ for both current and future detector configurations and variations in the detector’s performance due to bioluminescence activity remain below 1010% for most of the data-taking time. When implemented in KM3NeT’s realtime analysis system, the new, BDT-based, method led to a 4646% increase of the detection horizon. Even for the most pessimistic model considered here, the current detector configurations could probe the vast majority of CCSN candidates in the Milky Way and, once completed, KM3NeT will achieve full Galactic sensitivity. Hence, even with its current partial configuration, KM3NeT can play an important role in the detection and localization of the next Galactic supernova, in particular as part of the Supernova Early Warning System [al2021snews]. The results of this study highlight the versatile performance of the multi-PMT design to characterize environmental backgrounds and investigate low-energy phenomena at neutrino telescopes deployed in natural, dynamic environments.
@@ B.1 Signal simulation
Here, $\theta_{i}$ is the angle between the positron and neutrino directions. In this analysis, the neutrino direction is set to $\theta=\pi/3,$ and $\phi=4\pi/3$, corresponding to a CCSN located in the Galactic plane (most probable supernova location). $E_{p,i}$ is the positron energy, $E_{\nu,i}$ the neutrino energy and $\sigma$ the cross section of the process considered.
The CCSN simulation described above is used to assess the discriminating power of single-DOM observables, design a new single-DOM event selection strategy, and evaluate the sensitivity of KM3NeT to CCSNe. Simulated events are weighted according to the models used in the initial KM3NeT CCSN search [van2021km3net].
@@ B.2.1 Modeling radioactive decays
Radioactive decays are simulated using the KM3NeT software package OMGsim [colomer2017detailed]. This GEANT4-based package models the propagation of the decay products as well as the production and propagation of the Cherenkov photons. As for the CCSN signal, interaction vertices are simulated as being uniformly distributed within a sphere around a single DOM. A detailed technical description of this DOM is used to simulate photon propagation within its material and the final photoelectron production in the PMT photocathodes. The five dominant sources of radioactivity background in KM3NeT (using the respective decay rates in the deep seawater given in [albert2018long]) are the decays of 40K and 238U in water, and of 40K, 238U and 232Th in the DOM glass. The relative contributions of these decays to the total radioactivity background are given in Table 2. For these elements, decay vertices are simulated within a radius of up to $1\text{\,}\mathrm{m}$ around the DOM. This radius has been shown to contain $99.5\%$ of the decays for events with multiplicity $4$ and above.

### v2

@@ 5.2 Systematic uncertainties
The main source of uncertainties on the background rates is the extrapolation of the fraction of background events surviving the triggered event veto, or “veto acceptance”, when considering future detector configurations. Radioactivity backgrounds are not affected by the triggered event veto, whereas atmospheric muon backgrounds are strongly suppressed in a manner that is dependent on the detector geometry [5]. Consequently, for a given event sample, the veto acceptance is determined by two quantities: (1) the relative fraction of radioactivity events in the sample before the veto, which is obtained through the data-driven method described in Section 4, and (2) the fraction of atmospheric muon events that are found to survive the veto, evaluated using simulations. Following the procedure from [5], uncertainties on the veto acceptance are estimated by comparing data and simulations for the ORCA15 and ARCA21 configurations for each value of the multiplicity and the BDT discriminant. To avoid underestimating the systematic uncertainties, simulations are corrected before the triggered event veto is applied. The veto acceptance as a function of the multiplicity is presented in Figure 7 for both ORCA and ARCA, showing that discrepancies between data and simulations and the resulting systematic uncertainties are well under control.
@@
Systematic uncertainties on the neutrino interaction cross section, water optical properties, and PMT characteristics are also considered both for the cut optimization and when computing sensitivities to CCSN models. The uncertainty values from [5] are used for the cut optimization, and these values are re-evaluated when computing the sensitivities.
@@ 6 KM3NeT sensitivity to CCSN neutrinos
A key objective of this analysis is to position KM3NeT as a strong contributor to the Supernova Early Warning System (SNEWS) [8], which combines the information received from multiple detectors to reliably identify a CCSN (with a false alarm rate of one per century), locate the supernova using triangulation, and notify telescopes for prompt electromagnetic follow-up. The procedure developed in this paper is designed to identify potential CCSN neutrino signals and issue rapid alerts to this global network of observatories and detectors, following the realtime analysis procedure described in [6]. In this section, the sensitivity of the new CCSN analysis based on BDTs, incorporating the selection steps and systematic uncertainty evaluation procedure described above, is estimated for different detector configurations and CCSN models. The number of selected single-DOM events is computed within a $0.5\text{\,}\mathrm{s}$ time window, corresponding to the high-rate neutrino emission phase predicted by most models, and compared to the background expectation. As in the previous analysis [5, 6], the selection steps are optimized to maximize the distance up to which a CCSN with a neutrino flux corresponding to the $11~M_{\odot}$ model from [27] can be detected at $5\sigma$.
@@ 6.1 Selection procedure
In the previous CCSN analysis [5], the optimal multiplicity range was defined using a simple analytical approximation of the signal detection significance, neglecting systematic uncertainties. In this analysis, the signal detection significance is evaluated using the Rolke method [24], which accounts for systematic uncertainties.
@@
Four detector configurations have been considered to optimize the BDT and multiplicity cuts for a CCSN spectrum corresponding to the $11~M_{\odot}$ model considered in [5]: ARCA29, ORCA24, ARCA230 and ORCA115. The former represent the first ARCA and ORCA configurations for which the method presented here was used to search for CCSNe in real time. The latter represent the final ARCA and ORCA configurations. In each case, the background is modeled using data from the ARCA21 and ORCA15 configurations, respectively, and the signal is simulated using the average PMT efficiencies associated with these detectors. Since the typical $f_{A}$ values for ORCA15 range between $97$ and $98\%$, the associated background rates have been rescaled to model the $f_{A}>99\%$ scenario. With this choice of detector configurations, for ARCA29 and ORCA24, background rates can be rescaled linearly to account for the increase in detector size, as the variation of the background rate per DOM between the different configurations is negligible. For ARCA230 and ORCA115, an additional step is required to account for the reduced veto acceptance for background muons resulting from the increase in detector size. For this, the background evaluation method described in [5] is used: the background rates are evaluated from data before applying the triggered event veto, rescaled to match the increase in the number of DOMs, and then multiplied by the veto acceptances obtained from simulations (shown in Figure 7).
@@
The numbers of selected CCSN candidates after the optimization procedure are shown in the top panel of Figure 8 for the ORCA24 and ARCA29 detector configurations and for ORCA115-ARCA230. The optimal multiplicity ranges for the ORCA24-ARCA29 and ORCA115-ARCA230 detector configurations are 6116-11 and 7107-10, respectively, close to the range found in [5]. As in the previous KM3NeT CCSN analysis, ORCA backgrounds are less dominant thanks to ORCA’s higher atmospheric muon identification capabilities at low energies. For the signal, the $11~M_{\odot}$, $27~M_{\odot}$, and $40~M_{\odot}$ models used in [5] are shown for a CCSN at $10\text{\,}\mathrm{kpc}$. For all three models, the signal is dominating over the background, indicating that even with a partial configuration, KM3NeT is sensitive to most CCSN candidates in the dense Galactic bulge. To assess the impact of the BDT selection, the signal and background distributions obtained before this selection are shown in the bottom panel of Figure 8. The background rates differ from [5] as they have been evaluated using the method described in Section B.2 but the resulting impact on KM3NeT’s detection horizon is negligible. Applying BDT selection criteria significantly improves the signal-to-background ratio, in particular by improving the muon rejection rate at ARCA.
@@
Figure 8: Numbers of selected CCSN candidates for signal and background using the BDT-based method presented in this article (top) and the method from [5] (bottom), as a function of the multiplicity for ORCA24-ARCA29 (left) and ORCA115-ARCA230 (right). The multiplicity range shown for each detector configuration and selection procedure is the one providing the optimal sensitivity, as discussed in Section 6.1. For the signal, the $11~M_{\odot}$, $27~M_{\odot}$, and $40~M_{\odot}$ models used in [5] are shown. The light-blue upward-pointing and dark-blue downward-pointing triangles represent the ORCA and ARCA backgrounds, respectively. The dark blue histogram is the sum of the two backgrounds.
@@ 6.2 Sensitivities to CCSN models
Table 1: Systematic uncertainties of the analysis. The third column indicates the resulting systematic uncertainties on the signal and background rates denoted as (S) and (B), respectively. The evaluation of the uncertainties on the muon veto acceptance and single-DOM signal event modeling is described in Section 5.2. These uncertainties are quoted here for the optimal BDT cuts corresponding to the recent and final detector configurations. The monthly variability of the instrumentation efficiency [6] has been evaluated over a period of 11 year. The variations and uncertainties on the water transparency, the finite generation volume and the inverse beta decay (IBD) and elastic scattering (ES) cross sections are taken from [5]. The impact of the cross-section uncertainties for neutrino charged-current interactions with 16O are neglected, as these processes account for less than $1\%$ of the detected events.
@@
    In [5], the source of the corresponding uncertainty was considered to be the fraction of active PMTs. “Instrumentation efficiency variability” is a more accurate terminology which was introduced in a subsequent study [6].
Sensitivities are evaluated for the CCSN signals in the case of the $11~M_{\odot}$, $27~M_{\odot}$, and $40~M_{\odot}$ progenitor mass models and are shown in Figure 9, as a function of the distance to the CCSN, for the ORCA24-ARCA29 (left) and ORCA115-ARCA230 (right) detector configurations. Introducing the BDT selection leads to a $25$ to $30\%$ increase of the $5\sigma$ distance horizon as compared to the one obtained with the method from [5]. For the $11~M_{\odot}$ CCSN model, associated with the lowest neutrino flux, KM3NeT’s recent partial configurations have been sensitive to CCSNe up to a distance of $13\text{\,}\mathrm{kpc}$ and the detector will achieve full Galactic sensitivity when completed. The current reach of KM3NeT allows probing at least $79\%$ of Galactic CCSN candidates [1], $46\%$ more than with the previous analysis strategy [5]. The results highlight both KM3NeT’s potential for characterizing MeV neutrino bursts and the interest of investigating the structure of single-DOM hit coincidences to study low-energy signals.
@@
Figure 9: Sensitivity, in units of $\sigma$, as a function of the CCSN distance to Earth for the $11~M_{\odot}$ [27], $27~M_{\odot}$ [27], and $40~M_{\odot}$ [29] models already considered in the previous KM3NeT CCSN search [5]. Left: sensitivities for the ORCA24-ARCA29 detector configuration. Right: projected sensitivities for the completed ORCA115-ARCA230 detector. The sensitivities obtained with the selection from  [5] are indicated by the dotted lines. For ORCA115-ARCA230, these detection horizons agree with the ones from [5] despite the different background estimation method.
Figure 10: Sensitivity, in units of $\sigma$, as a function of the CCSN distance to Earth for the $11~M_{\odot}$ [27], $27~M_{\odot}$ [27], and $40~M_{\odot}$ [29] models already considered in the previous KM3NeT CCSN search [5]. In this figure, the performances of ARCA and ORCA are evaluated separately. The sensitivities for the ORCA24 (solid lines) and ARCA29 (dashed lines) detector configurations are indicated in the left panel, and the projected sensitivities for the ORCA115 (solid lines) and ARCA230 (dashed lines) detectors are indicated in the right panel.
@@ 6.3 Impact of variations in data-taking conditions
The sensitivities presented above have been evaluated assuming a fraction of active PMTs $f_{A}>99\%$. While this value is now achieved for more than $99\%$ of ARCA’s data-taking period, $f_{A}$ can reach values as low as $50\%$ at ORCA, where bioluminescence levels are typically higher. To account for these variations, the parameterization of the background rate dependence on $f_{A}$ proposed in [6] is adapted to the BDT-based single-DOM event selection procedure. The rate of single-DOM background events as a function of $f_{A}$ is shown in Figure 11 for ORCA18. For more than $95\%$ of ORCA18’s data-taking period, background rates can be successfully predicted using a parameterization obtained for the ORCA6 configuration. This predictability enables a reliable estimation of the expected background level when searching for CCSNe in real time, even when the bioluminescence activity is high. For ORCA, during $68\%$ of the data-taking time, $f_{A}$ will range from $91\%$ to $99\%$, resulting in variation of ORCA’s CCSN detection horizon by only $8\%$. The variation of the combined ORCA+ARCA detection horizon is expected to be even smaller due to the lower bioluminescence activity (hence typically larger $f_{A}$) at ARCA. Hence, KM3NeT’s CCSN identification capabilities are robust even in the highly dynamic environment of the Mediterranean Sea.
@@ 7 Conclusions
In this paper, a novel approach is proposed to improve the sensitivity of KM3NeT to CCSN neutrinos, by selecting single-DOM hit clusters based on their topology. A multi-variate analysis using BDTs has demonstrated that most of the hit information in a single-DOM event can be encapsulated into a small set of observables. With these observables, the ability to identify CCSN neutrino signals and separate them from radioactivity and atmospheric muon signatures, thereby allowing the estimation of the corresponding systematic uncertainties, is enhanced. Consequently, KM3NeT’s CCSN detection horizon is extended by $30\%$ for both current and future detector configurations and variations in the detector’s performance due to bioluminescence activity remain below 1010% for most of the data-taking time. When implemented in KM3NeT’s realtime analysis system, the new, BDT-based, method led to a 4646% increase of the detection horizon. Even for the most pessimistic model considered here, the current detector configurations could probe the vast majority of CCSN candidates in the Milky Way and, once completed, KM3NeT will achieve full Galactic sensitivity. Hence, even with its current partial configuration, KM3NeT can play an important role in the detection and localization of the next Galactic supernova, in particular as part of the Supernova Early Warning System [8]. The results of this study highlight the versatile performance of the multi-PMT design to characterize environmental backgrounds and investigate low-energy phenomena at neutrino telescopes deployed in natural, dynamic environments.
@@ B.1 Signal simulation
Here, $E_{p,i}$ is the positron energy, $E_{\nu,i}$ the neutrino energy and $\sigma$ the cross section of the process considered. $\theta_{i}$ is the angle between the positron and neutrino directions. In this analysis, the neutrino direction is set to $\theta=\pi/3,$ and $\phi=4\pi/3$, corresponding to a CCSN located in the Galactic plane (most probable location for an observable supernova). The impact of the neutrino direction on the sensitivity of KM3NeT is however negligible.
The CCSN simulation described above is used to assess the discriminating power of single-DOM observables, design a new single-DOM event selection strategy, and evaluate the sensitivity of KM3NeT to CCSNe. Simulated events are weighted according to the models used in the initial KM3NeT CCSN search [5].
@@ B.2.1 Modeling radioactive decays
Radioactive decays are simulated using the KM3NeT software package OMGsim [17]. This GEANT4-based package models the propagation of the decay products as well as the production and propagation of the Cherenkov photons. As for the CCSN signal, interaction vertices are simulated as being uniformly distributed within a sphere around a single DOM. A detailed technical description of this DOM is used to simulate photon propagation within its material and the final photoelectron production in the PMT photocathodes. The five dominant sources of radioactivity background in KM3NeT (using the respective decay rates in the deep seawater given in [9]) are the decays of 40K and 238U in water, and of 40K, 238U and 232Th in the DOM glass. The relative contributions of these decays to the total radioactivity background are given in Table 2. For decays occurring in water, vertices are simulated within a radius of up to $1\text{\,}\mathrm{m}$ around the DOM. This radius has been shown to contain $99.5\%$ of the decays for events with multiplicity $4$ and above.